Break-Even Sales and Cost-Volume-Profit Graph

For the coming year, Bernardino Company anticipates a unit selling price of $144, a unit variable cost of $72, and fixed costs of $640,800.

Instructions:

1. Compute the anticipated break-even sales in units.
units

2. Compute the sales (units) required to realize operating income of $244,800.
units

3. Construct a cost-volume-profit graph on paper, assuming maximum sales of 17,800 units within the relevant range. From your chart, indicate whether each of the following sales levels would produce a profit, a loss, or break-even.

$1,800,000
$1,598,400
$1,281,600
$964,800
$763,200

4. Determine the probable operating income (loss) if sales total 14,200 units. If required, use the minus sign to indicate a loss.
$  
Answer #1

a) Break even unit = Fixed cost/Contribution margin per unit = 640800/(144-72) = 8900 Units

2) Required unit = (640800+244800)/72 = 12300 Units

3) Calculate following

$1,800,000 Profit
$1,598,400 Profit
$1,281,600 Break even

$964,800

763200

Loss

Loss

4) Net income (loss) = Sales -Expense = (14200*72)-640800 = 381600